Belinda's capital gains and losses for two consecutive years are provided here. What is the effect of the gains and losses on Belinda's taxable income for each year?

Please include the explanation, thank you!

Belinda’s Capital Gains and Losses for Two Consecutive Years

-

First Year

Second Year

Short-term capital gain

$0

$4,000

Long-term capital gain

13,000

6,000

Short-term capital loss

(3,000)

(15,000)

Long-term capital loss

(12,000)

0

Group of answer choices

First year: $10,000; Second year $ 7,000;

First year: $(2,000); Second year $ (3,000);

First year: $(3,000); Second year $ (3,000);

First year: $(2,000); Second year $ (5,000);
Answer #1

Option is: D ( First year: $(2000); Second year $(5000)

First year:

short term = $0 -3000 = -3000

long term = 13000-12000 = 1000

Short term net loss = 1000-3000 = -2000

Second year:

Short term = 4000-15000 = -11,000

long term = 6000-0 = 6000

Long term net loss = 6000-11000 = -5000

So the option is D First year $(2000); Second year $(5000).